OpenStudy (anonymous):

\[d(y(u,v))=\frac{\partial y}{\partial u}du+\frac{\partial y}{\partial v}dv\]Why?

OpenStudy (anonymous):

Or: why aren't there half signs in front of each term?

OpenStudy (anonymous):

Right now I see the equation as\[dy=dy+dy=2dy\]

hartnn (hartnn):

those dy's are not variables, its not like xy = xy/2+xy/2 there is product rule involved (fg)' = fg'+f'g

OpenStudy (anonymous):

I don't understand how to expand the product rule into a function of 2 variables.

OpenStudy (experimentx):

that is just like linear approximation on a function with 2 independent variable.

OpenStudy (anonymous):

\[\frac{dz(a)y(b)}{d?}=\frac{dz}{da}y+\frac{dy}{db}z\]

OpenStudy (anonymous):

What is the question mark, or is my premise incorrect?

OpenStudy (anonymous):

Is the concept in the question a little like the gradient of a 2-variable function?

OpenStudy (anonymous):

With du and dv infinitesimal unit vectors?

OpenStudy (experimentx):

you know Taylor expansion of two independent variables?

OpenStudy (anonymous):

No

OpenStudy (experimentx):

oh ... this relation works our if \( \phi \) is function of x, y, z ... but what you wrote is more fundamental than this one. \[ d\phi = \nabla \phi \cdot d\vec r \]

OpenStudy (experimentx):

you know linear approximation right?

OpenStudy (anonymous):

No, but google says the first 2 terms of a function's Taylor E = its linear approximation. Correct?

OpenStudy (anonymous):

Single variable, that is.

OpenStudy (experimentx):

\[ f(x) = f(a) + f'(a)(x-a) + ... \implies f(x) - f(a) = \Delta f = {df \over dx}\Delta x \]

OpenStudy (anonymous):

\[f'(a)x-f'(a)a=f(x)-2f(a)\] I'm probably being an idiot, but I don't see it.

OpenStudy (experimentx):

f(x) - f(a) = this is delta f ... x - a = this is delta x

OpenStudy (anonymous):

I understand that, but not the equalsarrow part

OpenStudy (experimentx):

\[ f(x) = f(a) + f'(a)(x-a) \\ f(x) - f(a) = f'(x) (x-a) \\ \Delta f = f'(x) \Delta x\]

OpenStudy (anonymous):

OK- I've got that (sorry, thought that equals arrow was 'equal to')

OpenStudy (experimentx):

all right ... when x->a delta f = df delta x = dx this is linear approximation for a single variable function. do you know how to make linear approx for multivariate function?

OpenStudy (anonymous):

No

OpenStudy (experimentx):

use the same concept as single variable function ... use steps and ignore the dy dx (quadratic term) ... i'll give you in short.

OpenStudy (experimentx):

yeah that's the one.

OpenStudy (experimentx):

put f(x,y) - f(a,b) = df x-a = dx y-b = dy

OpenStudy (anonymous):

Oh, I see.

OpenStudy (anonymous):

I understand your answer algebraically, but is gradient a good conceptual model to hold about it?

OpenStudy (experimentx):

consider single it variable x \[ f(a+dx, b+dy) = f(a, b+dy) + {\partial f (a, b+dy)\over \partial x}_{x=a}\; dx --- (1)\] consider it single variable in y \[ f(a, b+dy) = f(a, b) + { \partial f \over \partial y} _{y=b}dy ---- (2)\\ {\partial f (a, b+dy)\over \partial x} dx = {\partial f (a, b)\over \partial x}_{x=a}dx + {\partial \; f\over \partial x \partial y}_{a,b} dx dy ---- (3)\] put everything at (1) ... and drop the quadratic terms. you will get desired result.

OpenStudy (experimentx):

probably this is more fundamental concept than gradient.

OpenStudy (anonymous):

Is it a branch of calculus that gradient is a twig of, or are the concepts completely unrelated?

OpenStudy (experimentx):

nothing is unrelated.

OpenStudy (anonymous):

Call the function of intensity of relatedness between two mathematical concepts a and b h(a,b). Is\[\left \langle h(a,b) \right \rangle<h(gradient , \text{ 2 variable linear approximation})\]?

OpenStudy (anonymous):

Taking the modulus of the values first, of course.

OpenStudy (experimentx):

both are very useful technique ... you use gradient more on vector calculus .. you use this approximation everywhere. especially on physics.

OpenStudy (anonymous):

So it's not used in vector calculus? What are some examples? Thanks for your time, but I've got to go.

OpenStudy (experimentx):

it's used everywhere.